A rectangle has dimensions 8 inches by 66 feet. Express its area in terms of square feet
Answer by rapaljer(4671)

8 inches=8/12 = 2/3 foot.

Area = L*W
=66*2/3
= 44 square feet

Dr. Robert J. Rapalje, Retired